OpenStudy (hpfan101):

Determine the constants a and b such that \[\lim_{x \rightarrow 0}\frac{ \sqrt[3]{ax+b}-2 }{ x }=\frac{ 5 }{ 4 }\] (Without the use of L'Hospital's Rule)

OpenStudy (michele_laino):

hint: using Taylor's expansion, around \(x=0\), I get this: \[\Large f\left( x \right) = \frac{{\sqrt[3]{b} - 2}}{x} + \frac{a}{{3\sqrt[3]{{{b^2}}}}} + o\left( x \right)\] where \(\large o(x)\) is a quantity which goes to zero as \(x\). So can we hope to have the limit, if the two subsequent conditions hold: \[\Large \left\{ \begin{gathered} \sqrt[3]{b} - 2 = 0 \hfill \\ \hfill \\ \frac{a}{{3\sqrt[3]{{{b^2}}}}} = \frac{5}{4} \hfill \\ \end{gathered} \right.\]

OpenStudy (michele_laino):

oops.. we can hope...

OpenStudy (thomas5267):

\[ \begin{align*} &\phantom{{}={}}\frac{\sqrt[3]{ax+b}-2}{x}\\ &u=ax+b\\ &x=\frac{u-b}{a}\\ &=\frac{\sqrt[3]{u}-2}{\frac{u-b}{a}}\\ &=a\left(\frac{\sqrt[3]{u}-2}{u-b}\right) \end{align*} \] Can we evaluate this? \[ \lim_{u\to b}\frac{\sqrt[3]{u}-2}{u-b} \]

OpenStudy (thomas5267):

Is the limit finite? \(\displaystyle \lim_{x \rightarrow 0}\frac{ \sqrt[3]{ax+b}-2 }{ x }\) does not look finite to me since \(\sqrt[3]{ax+b}-2\) will not blow up to infinity but \(\frac{1}{x}\) will.

OpenStudy (thomas5267):

Taking x to infinite might make more sense.

OpenStudy (thomas5267):

\[ \begin{align*} &\phantom{{}={}}\lim_{u\to {\color{red}\infty}}\frac{\sqrt[3]{u}-2}{u-b}\\ &=\lim_{u\to {\infty}}\frac{u^{-2/3}-2/u}{1-b/u}\quad\text{Divide up and down by }u\\ \end{align*} \] Can this be evaluated?

OpenStudy (thomas5267):

http://www.wolframalpha.com/input/?i=Limit [%28%28ax%2Bb%29^%281%2F3%29-2%29%2Fx%2C+x-%3E0] It shows that the limit of x to 0 doesn't exist. If the link doesn't work, input this into Wolfram|Alpha. Limit[((ax+b)^(1/3)-2)/x, x->0]

ganeshie8 (ganeshie8):

\(\lim\limits_{x \rightarrow 0}\dfrac{ \sqrt[3]{ax+b}-2 }{ x }\\~\\ =\lim\limits_{x \rightarrow 0}\dfrac{ \sqrt[3]{ax+b}-\sqrt[3]{8} }{ x }\\~\\ =\lim\limits_{x \rightarrow 0}\dfrac{ax+b-8 }{ x( \sqrt[3]{(ax+b)^2}+ \sqrt[3]{ax+b}\sqrt[3]{8} + \sqrt[3]{8^2} ) }\\~\\ \) It's tricky, but its not so hard to convince ourselves that \(b=8\) and \(a=15\)... I'll try and come up with some reasoning why they must work..

OpenStudy (thomas5267):

How does this even work? This is indeed the correct answer as verified by Wolfram|Alpha. Why does it not blow up to infinity? http://www.wolframalpha.com/input/?i=Limit [%28%2815x%2B8%29^%281%2F3%29-2%29%2Fx%2C+x-%3E0] Limit[((15x+8)^(1/3)-2)/x, x->0]

OpenStudy (thomas5267):

I guess you could divide up and down by x?

ganeshie8 (ganeshie8):

Ahh that looks like a good idea since we don't have the luxury to use L'hospitals... \(\lim\limits_{x \rightarrow 0}\dfrac{ \sqrt[3]{ax+b}-2 }{ x }\\~\\ =\lim\limits_{x \rightarrow 0}\dfrac{ \sqrt[3]{ax+b}-\sqrt[3]{8} }{ x }\\~\\ =\lim\limits_{x \rightarrow 0}\dfrac{ax+b-8 }{ x( \sqrt[3]{(ax+b)^2}+ \sqrt[3]{ax+b}\sqrt[3]{8} + \sqrt[3]{8^2} ) }\\~\\ =\lim\limits_{x \rightarrow 0}\dfrac{a+(b-8)/x }{ \sqrt[3]{(ax+b)^2}+ \sqrt[3]{ax+b}\sqrt[3]{8} + \sqrt[3]{8^2} }\\~\\ = \text{some finite number} + \lim\limits_{x \rightarrow 0}\dfrac{b-8 }{ x(\text{some finite number}) }\\~\\ \) then i hope we may argue that the overall limit exists only if \(b=8\)

OpenStudy (thomas5267):

If b neq 8 then 1/x blows up to infinity. If b=8 then b-8=0 and the second term vanishes?

OpenStudy (thomas5267):

Now how to determine a?

ganeshie8 (ganeshie8):

we may simply plugin b=8, evaluate the limit and set it equal to 5/4

OpenStudy (freckles):

In order for the limit to exist top will need to approach 0 since the bottom is approaching 0. So \[\sqrt[3]{a(0)+b}-2=0 \\ \text{ gives us } b=8\] plug back in and then I would apply the substitution to evaluate the limit (pretending I don't already know the limit) \[u=\sqrt[3]{ax+b} \\ x \rightarrow 0 \text{ so }u \rightarrow 2 \\ \lim_{u \rightarrow 2} \frac{u-2}{\frac{u^3-8}{a}}\] u^3-8 is easy to factor from here it should be easy to come up with a

OpenStudy (hpfan101):

I'm kind of confused about which process would work out for this problem. Would the taylor expansion one work? Based on the work shown on this thread, it looks like we would get a=15 and b=8

OpenStudy (anonymous):

@hpfan101, HP as in Harry Potter?

OpenStudy (hpfan101):

Yeah

OpenStudy (freckles):

I think there have been a couple of ways mentioned @hpfan101

OpenStudy (freckles):

Forgive if I don't mention your way... I see three ways. Taylor Rationalizing the numerator A substitution

OpenStudy (kainui):

If you're not allowed to use L'Hopital's rule, you're generally not allowed to use Taylor series either @Michele_Laino

OpenStudy (hpfan101):

Oh, ok. I see the three ways now.

OpenStudy (hpfan101):

@Kainui Oh, so Taylor expansion isn't a method I could use for this problem?